17 November 2009 Lab Meeting
- Oakridge deadline is 1 Feb 2010.
- PAPERS Drafted before MicroTAS
‡ Flu R01:Integration
- Set up new INTEGRATION MEETING for before 11/25.
† HDA (Lead: Jaephil, Team:Sonali)
- Start planning R01 for Submission on 2/5/10. MM, JD, CMK.
- HDA chip fabrication training will be after new process settled (comeback from mTAS).
- Two methods, Hot embossing and cutter plotting will be combined to reduce whole process time.
- Dry reagent storage for RT and HDA
- Set-up meeting for 2nd HDA design - Target (Flu or Bacteria), include RT or not, Heat source, address evaporation loss, Reagent delivery and storage solution. Consider marginating flows and include RT?
- Paper resubmission.
† Sample Concentration (Lead: Jane, Team: Jaephil)
- New design sent to make Rubber mold: outline octagon problem. Revised design sent yesterday.
- Jane will train Cassidy once more on Wednesday and Thursday this week.
- MM 040808 A/PR/8/34 flu stock with PFU = 2X10^4 available to Jane to train Cassidy in ultracentrifuging all tubes
- of virus, pour out sucrose,pipette remaining with 20uL pipette tip and pool for plaque assay for supplying virus
- Jane working on the cell lysate control. - have many tubes now, done
- Set up COMSOL and StarCD, look for governing equations for simulation of evaporation. - ongoing
† SPE Column Optimization for RNA. (Lead: Sonali, Team: Hussam, Jessie, Brendan)
- pH the sample. Optimum 6.6. Make sure that we are there.
- Gold standard Qiagen Kit.??
- silica nanoparticles from Fluka.
- Report Results today on 2 color dyes, best channel size for 1X 700 nm silica DNA, RNA and 3X 700 nm silica DNA, RNA
- Did adhesive tape instead of bonding help with 3X chips?
- Abandon 3X experiments with current channel designs. If 3X data are encouraging, develop a new channel design for 3X
- What are the next experiments? If types of silica -- Hussam and Jessie describe types today and order by Th 11/19/09
- Video with two color dyes and [red(primer) then blue, green, water+tween]
- repeat experiment reported on 10/27 with 700nm Silica 3X.
- controls for these: empties, non-silica monolith and full silica monolith, std. recipe.
- Load DNA in a straw of 5,50,100ul SPE, quick test of hypothesis: more SPE = more DNA, To be repeated with Alex and Mark
- Look into modifying Alex's pneumatic setup for the PATH guys. Instead of building the straw array over here NOTE: Sean
- PLanning net silica per channel experiments. with 700nm, 3x possible. Deliver around 11/5/09.
- 3X silica sealing with adhesive on 4 different width channel: Tried four, no success. SPE washout, crack forming within the SPE.
- Noticed some amounts of SPE coming out even on the channels that seemed to work in the beginning, so probably not retaining 3X silica.
- Would like to do experiments with different volumes of wash buffer, different flow rates, is it ok?
† PCR - CMI (Lead: Qingqing)
- QQ will work on the initial integration steps of SPE + RT (reservoir)+ PCR.
- Fourth design was made and tested with PCR reagent(Flu assay).ATCC RNA from Jessie was used.
- it works!
- patient sample will be tested next.
- primer for toxin A.
- product size and concentration was verified by bio-analyzer.
- on chip PCR was optimized with higher BSA concentration(0.02%).It did not work.
- 40 cycle chip and 0.03% (BSA concentration) will be tested.
- primer for toxin B.
-Both two pairs of primers for toxin B works, with a ct value around 13.
-on chip PCR has been done, the result has not been tested by bio-analyzer.
- PCR2 Paper formatted for LOAC this week.
need to get back from Cathie
- CMK: PCR 1 draft. Analytical Chem. MCK is running more simulations.
‡C. diff Project (Cathie, Sonali, Satish Singh, Lisa J., His post doc )
- Sonali to train Lisa on SPE.
- They don't have syringe pumps and have not asked for training yet. I offered training here which was declined.
- QQ to run test PCR on chip with genomic DNA and Toxin B primers.
‡ Coulter Flu Fraunhofer Project (Lead: Sonali, Team: Sonali, Jessie, Cathie, CMI Folks, Qingqing)
- Meeting 9am Monday (every other week)
- Data on AgPath vs Invitrogen one step RT-PCR with plasmid copies and chip extracted RNA on Mon by JC
‡ Agilent Automated Sample Preparation (Lead: Alex)
- So far tested on SNAP2:
- B. subtilis gDNA
- B. subtilis cells
- Planned tests SNAP2:
- B. subtilis cells (higher concentrations)
- MDCK gDNA
- HotDog Sample
- NANO-C: Straw making with two plugs in progress, first lysis test with B.Subtillis cells planned for this week.
- PATH: straw holder for Blackbird is machined
‡ COBRA (Lead: Jaephil, Team: Cathie, Jane for virus only, PHO folks)
- New metal piece design set.
- Evap paper 1st draft by the end of this week
‡ Biointerfaces group (Lead: MinCheol, Team: Cathie, MCK, Wong and Meller folks)
- New experiments happening now.
- Cathie will submit paper inquiry.
‡ CIMIT- Sepsis (Lead:Cathie, Team:TBA)
- IRB approved.
- Cathie submitted the companion BU IRB form for exemption. Revised, still still waiting.
‡ PATH Grant (Lead:Cathie, Team:Frank, Sean, Mark, Jake Trueb (ME), Suma)
- Phone meeting with PATH this week. 10/13.
- Senior Project Proposals due in less than one week! (23 Nov.)
- Frank: Unfortunately MIA due to interviews. 14-day unstabilized storage PCR results to be analyzed very soon.
- Frank: PVA arrived. Experiments to commence shortly.
- Sean: Design has been submitted to Alex and awaiting the product. I can then get back to working on the optimization of the straws, ie the best monolith composition for extracting DNA/RNA onto the straws
- Sean met with Alex for an integration of straws on black machine. They sketched a design. Sean will make a professional drawing and create a required parts list. Alex will supervise the design. Before handing in the job to the shop floor, the design will be presented to Cathie.
- Mark to begin initial DNA testing with shorter time periods (1,3 days, etc.) to be completed this(?) week. PCR training completed. Ready to begin testing. Waiting for new straw fixture to be ready for testing (where will it be stationed?).